OpenStudy (anonymous):

A fossilized animal that was killed by a volcanic eruption has 6.25 percent of its original amount of carbon-14 remaining; this radioisotope has a half-life of 5730 years. How long ago did the volcano erupt? 34,380 years ago 22,920 years ago 17,190 years ago 9,168 years ago

OpenStudy (anonymous):

I'll give a medal to whoever gets the answer!

OpenStudy (anonymous):

@Cuanchi

OpenStudy (anonymous):

@Rushwr

OpenStudy (cuanchi):

after 1 half-life the remained amount of C14 will be 50% of de original= 1/2 after 2 half-life the remained amount of C14 will be 50% of the 50% of the original= 1/4 =25% after 3 half-life the remained amount of C14 will be 50% of the 25% of the original= 1/8 = 12.5% after 4 half-life the remained amount of C14 will be 50% of the 12.5% of the original= 1/16 = 6.25 % if 1 Half life is 5730 years 2 Half life is 2 x 5730 years 3 Half life is 3 x 5730 years